OpenStudy (anonymous):

Add (8x^2-3x+23)+(7x^2+9x-46)

OpenStudy (anonymous):

119x - 23

OpenStudy (gw2011):

Combining like terms you get the following: 8x^2+7x^2=15x^2 -3x+9x=6x 23-46=-23 Therefore, the answer is: 15x^2+6x-23